OpenStudy (e.cociuba):

Calculus Help! MEDAL AWARDED! 1.) If y=x^2e^x, then dy/dx=? A.) 2xe^x B.) x( x+2e^x) C.) xe^x(x+2) D.) 2x+e^x E.) 2x+e

OpenStudy (e.cociuba):

@sjerman1 I need some help here. U said u did calc. rite?

OpenStudy (phi):

do you know the product rule?

OpenStudy (e.cociuba):

Yes, it states that dy/dx is the derivative of the first function times the second plus the first function times the derivative of the second. But it's a bit confusing.

OpenStudy (anonymous):

yeah hold on, phi is brilliant when it comes to calc, by the way

OpenStudy (e.cociuba):

k :)

OpenStudy (phi):

let u = x^2 and v= e^x now find u v' + v u' where v' and u' mean take the derivative

OpenStudy (phi):

do you know d/dx e^x ?

OpenStudy (phi):

d/dx e^x is one of the easiest derivatives .

OpenStudy (e.cociuba):

So, dy/dx = dydx(x^2) * (e^x) + (x^2) * dy/dx(e^x) dy/dx (x^2) is 2x dy/dx (e^x) is e^x so you get: 2x* e^x + x^2 * e^x But do i factor out x*e^x from both terms?

OpenStudy (phi):

after you get the answer, you might have to tweak it to match one of the choices.

OpenStudy (anonymous):

if im not mistaken, @phi yes you can factor out just the e^x but that doesnt really do to much

OpenStudy (e.cociuba):

K, I got xe^x(2+x) Wat do ya'll think @sjerman1 and @phi ??

OpenStudy (phi):

you have the correct result. now it's algebra to change it to match one of the answers.

OpenStudy (e.cociuba):

Yes so it's B.) Correct?

OpenStudy (phi):

do you have bad eyes?

OpenStudy (anonymous):

i would say no

OpenStudy (anonymous):

look at the other choices

OpenStudy (phi):

K, I got xe^x(2+x) does that match B ?

OpenStudy (anonymous):

xe^x(2+x) matches one on the list of choices...

OpenStudy (e.cociuba):

Haha um no but I do have glasses to wear sometimes lol :)

OpenStudy (e.cociuba):

Yes it matches B... Thanks guys for the help. I appericate it. Thanks @sjerman1 and @phi :)

OpenStudy (anonymous):

nooooooo. look at the other choices...

OpenStudy (phi):

If you think it matches B you should look more carefully.

OpenStudy (anonymous):

xe^x(2+x) A.) 2xe^x B.) x( x+2e^x) C.) xe^x(x+2) D.) 2x+e^x E.) 2x+e

OpenStudy (e.cociuba):

Opps my badd I meant to say C.) srry (typing error)

OpenStudy (anonymous):

remember since (2+x) both 2 and x are positive it doesnt matter which order you write them in -> (2+x) or (x+2)

OpenStudy (anonymous):

Good job!

OpenStudy (e.cociuba):

Haha Thanks :) Thanks for the help ;)